How do we solve y''+y=cos x by variation of parameters method?
1
Expert's answer
2021-09-02T07:25:41-0400

The homogeneous differential equation


"y''+y=0"

Characteristic (auxiliary) equation


"r^2+1=0"

"r=\pm i"

The general solution of the homogeneous differential equation is


"y_h=C_1\cos x+C_2\sin x"




"W(y_1, y_2)=\begin{vmatrix}
 y_1 & y_2 \\
 y_1' & y_2'
\end{vmatrix}=\begin{vmatrix}
 \cos x & \sin x \\
 -\sin x & \cos x
\end{vmatrix}"


"=\cos^2x+\sin^2x=1"

"W_1=\begin{vmatrix}
 0 & y_2 \\
 \cos x& y_2'
\end{vmatrix}=\begin{vmatrix}
 0 & \sin x \\
 \cos x & \cos x
\end{vmatrix}=-\sin x\cos x"


"W_2=\begin{vmatrix}
 y_1 & 0 \\
 y_1' & \cos x
\end{vmatrix}=\begin{vmatrix}
 \cos x & 0 \\
 -\sin x & \cos x
\end{vmatrix}=\cos^2 x"

"C_1'=\dfrac{W_1}{W(y_1, y_2)}=\dfrac{-\sin x\cos x}{1}=-\sin x\cos x"

"C_1=\int(-\sin x\cos x)dx=-\dfrac{1}{2}\sin^2x+C_3"

"C_2'=\dfrac{W_2}{W(y_1, y_2)}=\dfrac{\cos^2 x}{1}=\cos^2x"


"C_2=\int(\cos^2 x)dx=\dfrac{1}{2}\int(1+\cos(2 x))dx"

"=\dfrac{1}{2}x+\dfrac{1}{4}\sin(2x)+C_4=\dfrac{1}{2}x+\dfrac{1}{2}\sin x\cos x+C_4"


"y=(-\dfrac{1}{2}\sin^2x+C_3)\cos x"

"+(\dfrac{1}{2}x+\dfrac{1}{2}\sin x\cos x+C_4)\sin x"

The general solution of the given differential equation is


"y=C_3\cos x+C_4\sin x+\dfrac{x\sin x}{2}"